OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. -15, -6, 3, 12, ...

OpenStudy (john_es):

Arithmetic sequences are defined by \[a_n=a_0+d(n-1)\] where d is the distance between two consecutive terms, and a0 is the first term of the sequence. In this case, you have d=12-3=9, and a0=-15, so, \[a_n=-15+9(n-1)=9n-24\]